Fermi National Laboratory
 

Operation Summary for August 28, 2009 - August 31, 2009

 

Friday August 28

Shutdown.

During the day shift, operators prepared the Main Injector (MI) NuMI target area for a morning access. Later in the afternoon, Operators interlocked the Antiproton Source (Pbar) area and turned on a power supply (D:H717).

During the evening shift, D:H717 tripped off on a temperature indication. Operators found the magnet’s LCW valve wasn’t all the way open; They opened it. Per instructions, Operations turned on the Pbar lithium lens (D:LNV).

Saturday August 29

Shutdown.

During the midnight shift, operators checked on the Tevatron sector E3 temporary sump pump and found that FESS personnel had installed a new permanent pump. Operators closed the MI distributed ground knife switches.

During the day shift, operators checked all the MI correction element power supplies to make sure they were following their ramps; they were. Operators search and secured F-sector and MI. Vacuum technicians reported finding a vacuum leak on a beam pipe downstream of the MI-40 Lambertson.

During the evening shift, operators completed a MI hipot. Operators search and secured MINOS. Operators turned on the MI critical devices. Late in the shift, interlock techs began interlock-testing Pbar.

Sunday August 30

Shutdown.

During the midnight shift, Operations reported a problem with the Pbar beam stops.

During the day shift, Operations reported that the lithium lens had tripped off along with the sweeper magnet power supplies; a Pbar expert helped operators reset all the devices.

During the evening shift, a Klystron RF station (KRF3) tripped off for the first of three times over the next six and a half hours. It reset.

Monday August 31

Shutdown.

During the midnight shift, Operations prepared MI for access.

The Plans for Monday and for the rest of the Week

Shutdown work continues. A new MI beam pipe will be installed today and then begin a five-day bake. The new CUB chiller #3 will have a nitrogen leak test today. Linac is on and Booster is at five turns. MiniBooNE should begin taking beam on Tuesday, 9/1/09.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 09/04/2009    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory